ys 70 eoni idence interval for μο 4. The following table contains the minimum wage and unemployment rates for a sample of eight states. State 2018 Unemployment | Minimum | Rate (96) Wage Alabama Arizona Arkansas S7 $11 $8 California $11 4 Colorado $10 Connecticut $10 $8 $13 Delaware DC 4 Graph unemployment and minimum wage, with unemployment on the vertical axis. Draw a line that best fits the points and label the intercept and slope. You can use your answer from question 4.c for these labels. a) b) Compute the sample variance of unemployment and minimum wage and compute their covariance. Estimate the regression of a states unemployment rate on its minimum wage and write the resulting regression equation c) unempi = At Aminteagei d) e) Interpret the coefficient on minimum wage in a sentence. How would your answers to 4.c and 4.d change if the unemployment rate was expressed as decimals instead of percentages?
Answer #1

Answer a)

1-5 1.0 Minimum wage 0 0uneuploymeut Mth Wage Rate 49 98 12 55 64 94 121 44 1oo 4 lo 64 32 69 18 4 0 Y 33 -4.125 x-Y-9H5 331-48x33 = 331-391。45 8ー ·488-460-5 ー 488-4828 1-5 Nous flo= 4-125_-336 4x4.45 = .8151 여 = 3.989 Now Covasiauce oate -3364 % unita Anw